What are the nominal and effective costs of trade credit under the credit terms of 3/15, net 30? Assume a 365-day year. Do not round intermediate calculations. Round your answers to two decimal places.

Nominal cost of trade credit:   % ?

Effective cost of trade credit:   % ?
Answer #1

Nominal Cost of Trade Credit = [Disc.%/ (100 - Disc.%)] x

[365/ (Days Credit Outstanding - Discount Period)]

= [3/(100 - 3)] x [365 / (30 - 15)] = 3.09% x 24.33 = 75.26%

Effective Cost of Trade Credit = [(1+(Disc.%/100 - Disc.%)][365/(Days Credit Outstanding - Discount Period)] - 1

= [1.0309]24.33 - 1 = 2.098398 - 1 = 1.098398, or 109.84%